Heavy Metal Corporation is expected to generate the following free cash flows over the next five​ years:

Year

1

2

3

4

5

FCF​ ($ million)

52.3

67.2

76.4

74.4

82.2

After​ that, the free cash flows are expected to grow at the industry average of 4.2% per year. Using the discounted free cash flow model and a weighted average cost of capital of 14.5%​:

a. Estimate the enterprise value of Heavy Metal.

b. If Heavy Metal has no excess​ cash, debt of $291 ​million, and 40 million shares​ outstanding, estimate its share price.
Answer #1

a)

value of firm = Present value of FCF + Horizontal value

Horizontal value = FCF next year/(Required return - growth rate)

=>

horizontal value = 82.2 * 1.042/(0.145-0.042)

= 831.576699029 million

enterprise value of firm = 52.3/1.145 + 67.2/1.145^2 + 76.4/1.145^3 + 74.4/1.145^4 + 82.2/1.145^5 + 831.576699029/1.145^5

= 655.43 million

b)

total value = 655.43 - 291 million

stock price = 655.43 - 291/40

= 9.11